# CleAR: Robust Context-Guided Generative Lighting Estimation for Mobile Augmented Reality

### v1

@@ 6. Evaluation
We evaluate the performance CleAR using an AR application testbed, a set of data-driven benchmarks, and a user study. The testbed-based evaluation shows the real-world lighting estimation quality of CleAR. It also allows us to profile and break down the system latency. The data-driven benchmarks evaluate the lighting estimation quality of CleAR on several test datasets. To evaluate the robustness of CleAR’s lighting estimation, we introduce a new robustness testing protocol and a robustness testing dataset generated from the Laval dataset. The user study explores CleAR lighting estimation quality from the perspective of human perception.
@@ 6.1. Testbed-Based Evaluation
We implement a Unity-based cross-platform AR application to demonstrate the lighting estimation capability of CleAR. Our application implements an object placement action that allows AR users to visualize 3D objects in their physical environment at selected locations. This demonstration application mimics the behaviors of virtual object staging applications and shows potential improvements in application visual quality. When using the application, users are requested to select an environment location to pace objects based on on-screen visual instructions. After the object placement action, we request lighting estimation from CleAR and apply the result environment map for object rendering.
Figure 8. AR object rendering results. We show AR virtual object rendering qualitative comparison to other environment lighting acquired by unwrapping a physical mirror ball (Debevec, 2006) (reference), ARKit (Apple, 2022) (commercial), and LitAR (Zhao et al., 2022) (academic). Compared to ARKit, CleAR estimates the environment map with significantly more realistic visual details. Compared to LitAR, CleAR can generate more visually coherent far-field environment map details, while LitAR can only generate blurry ones.
6.1.2. Qualitative Visual Comparison.
We implement a Unity-based cross-platform AR application to demonstrate the lighting estimation capability of CleAR. Our application implements an object placement action that allows AR users to visualize 3D objects in their physical environment at selected locations. This demonstration application shows the potential impact of photorealistic rendering on immersive AR applications, such as AR shopping (Jiang et al., 2021). After the object placement action, we request lighting estimation from the testing lighting estimation system and apply the result environment map for object rendering in Unity. We followed the LitAR (Zhao et al., 2022) experiment setup to compare lighting estimation system performance under three different scenes. In Figure 8, we compare between environment lighting acquired by unwrapping a physical mirror ball (Debevec, 2006), LitAR (Zhao et al., 2022), and ARKit (Apple, 2022). The mirror ball captured environment map represents a reference physical environment lighting at the estimation position. To compare fairly to LitAR, we used the same near-field dense environment reconstruction data as our lighting estimation input. Overall, we observe that CleAR outputs the best environment map image quality among estimated environment maps (ARKit and LitAR). CleAR generates environment maps with more visual details than LitAR and fewer artifacts than ARKit. Particularly, CleAR can generate detailed far-field environments while LitAR can only generate blurry ones. The improved environment map quality also helps more visually coherent virtual object rendering.
6.1.3. System Performance Breakdown
We conduct a comprehensive system profiling to quantify the system runtime latency. We measure the end-to-end system latency of CleAR and other methods on a high-end workstation PC with an RTX 4090 GPU. In Table 1, we break down the end-to-end system latency measurement results. In the default configuration, CleAR generates five LDR completion results. Overall, the end-to-end inference takes 3269ms, mostly consisting of the generative LDR completion time. When combined with semantics in LDR completion, our systems take 3269ms to generate five LDR environment map completion results. We observe that the main impact factor for the LDR completion process is the number of environment map outputs. We notice that increasing generation output introduces a near-linear yet slow increase in the estimation latency. Although our default system configuration uses five outputs, the estimation latency is only about double the time of using a single estimation output.
Table 1. System performance breakdown. We show the breakdown of system latency under different generative estimation settings for CleAR. Our system can finish end-to-end environment map generation as fast as 2383ms. Under the default configuration, our system uses partial environment RGB in the inputs and generates five environment map completion results. The default estimation latency is 3269ms.
@@ 6.1.3. System Performance Breakdown
Data preparation 12 ($\pm$ 1.2)
LDR completion x1 (RGB) 1130 ($\pm$ 11.4)
LDR completion x3 (RGB) 1574 ($\pm$ 17.6)
LDR completion x5 (RGB) 2012 ($\pm$ 19.5)
LDR completion x7 (RGB) 2492 ($\pm$ 18.1)
Output processing 64 ($\pm$ 9.2)
HDR estimation x5 1166 ($\pm$ 10.7)
End-to-end System Avg. Time (ms)
@@ 6.2. Data-Driven Evaluation
6.2.1. Robustness Evaluation Data Generation.
Evaluating the robustness of lighting estimation systems is particularly challenging because it requires controlled lighting condition changes. To avoid costly real-time data capturing, we propose an image editing-based method that creates variants of standard lighting estimation testing datasets to represent diverse lighting conditions while maintaining the original environment map visual context. To do so, we use an informed augmentation selection-based method with the following three steps. First, we generate a set of edited variants of the Laval dataset by applying a uniform scaling term $s$ to all environment map images. Then, we measure the total light intensity and average color temperatures using the measurement method introduced in §3. Finally, we sample a range of edited dataset variants by comparing the newly measured light intensity and color temperature values to the original values. The scaling term $s$ are selected from $[0.25,4]$ using a step of $0.125$. For the light-intensity editing, the scaling term is uniformly applied to all three color channels, while only red and blue channels are scaled with $s$ and $1/s$ for color temperature editing (Afifi and Brown, 2020). Figure 9 shows examples of the edited environment maps. We created a set of augmented Laval indoor datasets using this method with edited average light intensity and color temperatures in $[-20\%,20\%]$ and $[-50\%,50\%]$.
Figure 9. Experiment protocol. Our experiment includes two protocols. For the three-sphere protocol (left), we use environment maps (middle) from the Laval dataset to render virtual spheres of different material properties. The rendered spheres provide us a proxy to understand the impact of lighting estimation results on virtual object rendering. We augment the Laval dataset to create diverse training and test data in different lighting conditions. We use the robustness testing protocol (right) to understand the robustness of lighting estimation results. We generate the robustness testing data by editing the Laval dataset environment maps’ ambient light intensity and color temperatures. In the example, we augment an environment map (middle) with four variants with different light intensities and color temperatures. The results (right) of light intensity augmentation are marked as dark and bright, and the results of color temperature augmentation are marked as warm and cold.
@@
To quantitatively evaluate the lighting estimation quality, we use several datasets and protocols to evaluate environment map estimation accuracy, diversity, and robustness. We use the original Laval dataset and our generated robustness testing datasets for the evaluation. We test our system against recent state-of-the-art lighting estimation models, including DiffusionLight (Phongthawee et al., 2023), StyleLight (Wang et al., 2022), and OmniDreamer (Akimoto et al., 2022). Among these baselines, DiffusionLight is the most similar to our method design because it uses a diffusion model-based pipeline for generating environment maps. Our experiment includes standard environment map evaluation on the Laval testing dataset. We also use two specialized experiment protocols to understand the robustness of our lighting estimation results and the impact of lighting estimation on virtual object rendering. Specifically, we use the following two protocols to conduct our evaluation:
@@
The three-sphere protocol. The three-sphere evaluation protocol (Wang et al., 2022; Phongthawee et al., 2023) is a proxy method to evaluate the HDR lighting estimation accuracy. The method evaluates the lighting estimation accuracy through pixel-wise error measurements on three rendered spheres with different material properties. Specifically, the used sphere materials are: matte, silver matte, and silver. Figure 9 shows a set of rendered virtual spheres used in this evaluation protocol. Each material represents a mixture of specific reflectance and roughness properties. Compared to direct environment map-wise comparison, this evaluation protocol helps us understand the impact of lighting estimation results when applied to object rendering.
@@
The robustness testing protocol. We design an experiment protocol to test the generalizability and robustness of lighting estimation systems under different lighting conditions, specifically with different light intensity and color temperatures. In this protocol, we test lighting estimation systems on the previously generated robustness evaluation dataset. Then, we compare estimated environment maps with the ground truth and calculate pixel-wise errors. This protocol directly calculates pixel-wise errors on the LDR image domain because it focuses on the lighting condition-related color appearance differences between environment map images.
@@
Evaluation Metrics. To quantify the quality of the generative lighting estimation pipeline, we evaluate the following aspects: (i) diversity of generated environment maps, and (ii) environment map estimation accuracy. Following previous research works (Wang et al., 2022; Akimoto et al., 2022; Phongthawee et al., 2023; Gardner et al., 2019), we use the FID score (Seitzer, 2020) to measure the diversity of environment generation. For accuracy evaluation, we use three scale-invariant image-based metrics to evaluate lighting estimation accuracy on rendered virtual spheres: scale-invariant Root Mean Square Error (si-RMSE) (Eigen et al., 2014), Angular Error (LeGendre et al., 2019), and normalized RMSE. For the last metric, we follow (Phongthawee et al., 2023) to map the 0.1st and 99.9th value percentiles to 0 and 1. We apply these evaluation metrics to measure the lighting estimation quality on the LDR environment completion results and the final HDR environment map output.
Figure 10. Qualitative comparisons. We show examples of lighting estimation results on three scenes and the corresponding rendered virtual sphere images. On the left, we show the input camera view for lighting estimation systems. On the right, we show rendered virtual spheres (row top) and estimated environment maps (row bottom).
a Accuracy
b Latency
Figure 11. End-to-end qualitative comparison. We compare the end-to-end estimation accuracy (11a) and estimation latency (11b) between CleAR and two state-of-the-art works (Wang et al., 2022; Phongthawee et al., 2023). Our method achieves the lowest estimation error on all three object material types in the three-sphere evaluation protocol. This indicates the lighting estimation results from our system can provide the best visual quality support for virtual object rendering. For estimation latency, our method significantly outperforms other generative model-based methods.
6.2.3. End-to-End Estimation.
In Figure 10, we show qualitative comparisons between CleAR and competing methods in three scenes, representing the neutral, warm, and cool lighting conditions. For each set of comparisons in the figure, we include the input camera view, estimated environment map, and rendered virtual spheres overlaid on the original camera view. The rendered virtual spheres are also examples of the three-sphere evaluation protocol. Compared to other methods, CleAR allows more accurate overall color tones rendering and creates more diverse reflection details on the rendered virtual spheres. The high-quality environment visual structure details contribute to the improved reflection details.
Figure 11 highlights the quantitative comparison between CleAR and two SoTA methods (Phongthawee et al., 2023; Wang et al., 2022). Overall, CleAR achieves leading estimation accuracy than comparing methods in terms of scale-invariant RMSE on the three-sphere protocol evaluation. In Figure 11b, we show that CleAR achieves lower lighting estimation latency than state-of-the-art methods by large margins. In particular, compared to a diffusion model-based method (DiffusionLight (Phongthawee et al., 2023)), CleAR takes significantly less time (110X) This reduction in estimation latency is due to CleAR’s fewer model inferences and lower per-inference latency compared to DiffusionLight. The estimation latency reduction allows faster virtual object rendering updates and more responsive user experiences in mobile AR applications.
Table 2. Ablation evaluation on the three-sphere protocol. On the standard testing settings with single 75 camera view input (row 1 - 3), our system achieves leading accuracy when compared to StyleLight (Wang et al., 2022) and DiffusionLight (Phongthawee et al., 2023) on the three-sphere evaluation protocol. More camera observations and environmental semantics will significantly reduce our system’s estimation error.
Scale-invariant RMSE $\downarrow$ Angular Error $\downarrow$ Normalized RMSE $\downarrow$
Method Diffuse Matte Mirror Diffuse Matte Mirror Diffuse Matte Mirror
StyleLight (Wang et al., 2022) 0.13 0.31 0.55 4.24 4.74 6.78 0.23 0.40 0.51
DiffusionLight (Phongthawee et al., 2023) 0.14 0.33 0.60 2.14 3.42 5.94 0.20 0.36 0.43
Ours w. RGB+semantics 75 FoV 1 view 0.06 0.14 0.28 4.41 5.91 8.89 0.17 0.20 0.37
Ours w. RGB+semantics 110 FoV 1 view 0.05 0.13 0.27 4.19 5.77 8.66 0.16 0.19 0.35
Ours w. RGB+semantics 110 FoV 3 views 0.03 0.10 0.24 3.89 5.21 8.23 0.13 0.17 0.31
Ours w. RGB+semantics 110 FoV 5 views 0.02 0.09 0.22 3.71 5.13 8.17 0.11 0.16 0.29
Ours w. RGB+full semantics 110 FoV 3 views 0.02 0.07 0.20 3.59 4.96 8.09 0.10 0.15 0.27
6.2.4. Context Data Impact Ablation.
@@
In Table 2, we show more comparisons between CleAR and two state-of-the-art method (Wang et al., 2022; Phongthawee et al., 2023) using the three-sphere evaluation protocol. We follow the setups in (Wang et al., 2022; Phongthawee et al., 2023) to use a single centered view of a 75 horizontal FoV camera as input. Under the same environmental observations, CleAR achieves lower scale-invariant and normalized RMSE values across all three virtual sphere types. Noticeably, for the mirror sphere, the sphere with the most challenging material, CleAR achieves approximately $50\%$ reduction in the measured scale-invariant RMSE values. On the other hand, we noticed that CleAR generates environment maps with slightly higher angular errors. Upon further inspection, we suspect that the diverse environment map image details potentially caused the higher angular error because the angular error metric is sensitive to pixel-wise differences.
@@
This setup configuration represents a baseline performance of CleAR because our system can take in more AR context data during the runtime. Therefore, we conducted further comprehensive evaluations on several configurations over camera FoVs, the number of input views, and the usage of semantic maps. Note that the ambient light data has been included in all tests in Table 2 because it is inseparable from our estimation pipeline design. The added environment observations through new views and increased FoVs can both improve the accuracy of lighting estimation. Additionally, full-scene semantic maps are important guiding information for high-accuracy lighting estimation. Full-scene semantic maps can often be obtained as floor maps or scene design layouts in real-world applications. Our testing result suggests that sharing scene semantics can greatly benefit lighting estimation accuracy with our system.
@@
6.2.5. Robustness Evaluation.
We test the lighting estimation robustness of CleAR under several environmental lighting conditions with the robustness testing protocol. For comparison, we test our system against two recent lighting estimation methods that are built on top of generative models: OmniDreamer (Akimoto et al., 2022) and StyleLight (Wang et al., 2022)  [^6] We were unable to complete the robustness testing for DiffusionLight (Phongthawee et al., 2023) within a reasonable time given its high inference cost and the large scale of data needed.. The lighting estimation error rates are calculated as normalized RMSE values on estimated LDR environment maps. In Figure 12, we observe that the accuracy of both OmniDreamer and StyleLight are affected by the changing lighting intensities and color temperatures. Particularly, the lower environmental lighting intensity significantly increases the lighting estimation error rates on existing models. The observed error increases potentially are caused by the unbalanced training datasets used by these models. While CleAR, on the other hand, shows consistently lower estimation error rates on different light intensity and color temperature conditions.
6.2.6. Evaluation of LDR Environment Map Completion
For the completion of the LDR environment map, we specifically examine the pixel-wise errors and the overall diversity of the environment map content. The former assesses CleAR’s LDR lighting estimation accuracy, and the latter evaluates the generation richness of CleAR generative estimation pipeline. When assessing the accuracy of the LDR estimation, we also include adaptive generation post-processing and candidate selection operations on the raw LDR environment map completion model output. Following (Akimoto et al., 2022; Wang et al., 2022), we adopt two methods for measuring the generation content diversity: (i) calculate the FID score on the full estimated LDR environment map, and (ii) converting the environment map into a cube map and calculate FID scores on each face without the top and bottom faces as these two faces containing little information. In Table 4, we show that our LDR environment map completion model can output environment map images with greater diversity than state-of-the-art models. Our LDR completion model outperforms other models by 27% to 370% in the first calculation method and 4% to 344% in the second calculation method. This observation confirms the effectiveness of our large LDR environment map completion dataset. Next, we compare the accuracy of the LDR environment map completion. Under the same environment observations as (Wang et al., 2022; Phongthawee et al., 2023), our method achieves comparable results to DiffusionLight even though our LDR completion model is based on a smaller-sized pre-trained diffusion model.
Table 3. Environment map generation diversity.
@@ 6.2.6. Evaluation of LDR Environment Map Completion
Table 4. LDR environment map estimation accuracy.
Method RMSE $\downarrow$ Si-RMSE $\downarrow$
Angular
Error $\downarrow$
StyleLight (Wang et al., 2022) 0.246 0.271 5.814
DiffusionLight (Phongthawee et al., 2023) 0.187 0.303 4.412
Ours 0.195 0.262 6.762
a Color Appearance Matching
b Output Selection
@@
Figure 13. Analysis of estimation refinement. (13a) Our color appearance matching technique can effectively improve overall estimation accuracy. (13b) We compare our output selection method to two straw-man baselines: single estimation output and random selection. Our method reduces the estimation of RMSE and its standard deviation compared to other baselines.
6.2.7. Analysis of Estimation Refinement
@@
We evaluate the performance of our on-device estimation refinement components on the impacts of lighting estimation accuracy. In Figure 13a, we show that our color appearance matching technique improves the overall estimation accuracy by 31%. The estimation refinement technique allows our system to achieve high-quality estimation with limited estimation outputs. By default, our system only requires five generation outputs while DiffusionLight (Yang et al., 2023) requires more than 90 generation outputs. Reducing the required generation output will also reduce the estimation latency and computation resources requirements during deployment. Next, we evaluate the effectiveness of our generation output selection policy. In Figure 13b, we show that our color-matching technique can reduce the estimation error rate compared to other selection methods. We observed that, with our output selection policy and a total of five generation outputs, CleAR can reduce 15% of the estimation error compared to using a single generation output. Furthermore, when using five generation outputs, our generation output selection component can reduce the average estimation error rate by 11% compared to random selection.
7. User Study

### v2

@@ 6. Evaluation
To comprehensively evaluate CleAR, we use a combination of quantitative and qualitative methods to assess key aspects of system performance, including generation latency, output quality, and user perception. We conduct a lab-based evaluation using an example AR application built on top of CleAR to assess rendering quality and latency. Our data-driven evaluation compares CleAR to several state-of-the-art lighting estimation methods (Phongthawee et al., 2023; Wang et al., 2022; Akimoto et al., 2022) using both the standard three-sphere protocol and our proposed robustness testing protocol under varied lighting conditions. In addition, we use an online survey to understand the impact of lighting estimations on human perceptual preferences. Our results demonstrate that CleAR achieves high-quality lighting estimation with low latency. For example, our end-to-end estimation results show that CleAR achieves the best performance in two image-based metrics while taking significantly less time. Participants also rated CleAR-generated lighting as producing better renderings for most virtual objects.
@@ 6.1. Testbed-Based Evaluation
We conduct our evaluation using an example AR application running on a 4th-generation 11-inch iPad Pro with a built-in LiDAR sensor. The edge server is a high-end workstation with an Intel Core i9-13900K CPU, 128GB of RAM, and an NVIDIA RTX 4090 GPU. The edge server runs an Ubuntu 22.04 operating system. The mobile device and edge server are connected via an 802.11ac Wi-Fi network. To ensure consistent and reliable results, all system runtime measurements are performed 10 times, and the reported values are the average across these runs.
Figure 7. AR object rendering results. We show AR virtual object rendering qualitative comparison to other environment lighting acquired by unwrapping a physical mirror ball (Debevec, 2006) (reference), ARKit (Apple, 2022) (commercial), and LitAR (Zhao et al., 2022a) (academic). Compared to ARKit, CleAR estimates the environment map with significantly more realistic visual details. Compared to LitAR, CleAR can generate more visually coherent far-field environment map details, while LitAR can only generate blurry ones.
Table 1. System performance breakdown. (Left) We show the breakdown of system latency under the default generative estimation settings, which use five LDR completion outputs and do not use semantic maps in the input. In the left table, on-device operations are marked in gray, edge operations are marked in blue, and network operations are marked in green. (Right) We show the end-to-end system execution time (excluding network time) under various system configurations. Our system can generate an environment map as fast as 2,383 ms. Under the default system configuration, which is marked in bold font, the end-to-end execution time is 3,269 ms.
@@
Device data preparation 0.5 ($\pm$ 0.1)
AR Data Offloading 31 ($\pm$ 10.1)
LDR completion x5 1957 ($\pm$ 9.3)
LDR env. map retrieval 72 ($\pm$ 8.1)
Estimation refinement 3 ($\pm$ 0.2)
Refinement result synchronization 10 ($\pm$ 0.6)
High-intensity map estimation 1166 ($\pm$ 10.7)
High-intensity map retrieval 30 ($\pm$ 8.9)
System Configuration Avg. Time (ms)
@@
6.1.2. Qualitative Visual Comparison.
We showcase CleAR’s lighting estimation quality with our testbed AR application in real-world environments. We followed the experiment setup of LitAR (Zhao et al., 2022a), a recent lighting estimation mobile system, to compare the quality of lighting estimation on rendered virtual sphere images. In Figure 7, we show the visual comparison between two baselines: LitAR (Zhao et al., 2022a) and ARKit (Apple, 2022). We also include a physical mirror ball as a reference for physical environment lighting. For fair comparison, we use the same near-field reconstruction data from LitAR as the RGB input to CleAR. Overall, we observe that CleAR outputs the best environment map image quality among estimated environment maps (ARKit and LitAR). CleAR generates environment maps with more visual details than LitAR and fewer artifacts than ARKit. Particularly, CleAR can generate detailed far-field environments while LitAR can only generate blurry ones. When comparing rendered virtual mirror balls to the physical reference mirror ball, CleAR produces the closest visual match.
6.1.3. System Performance Breakdown
Table 1 shows CleAR’s component-wise latency and its end-to-end execution latency under different configurations. In the default configuration, CleAR takes RGB images and ambient light data from the AR device as input to generate five LDR completion variants and one corresponding high-intensity map. The end-to-end execution takes 3,269 ms, primarily consisting of the generative LDR completion time. When combined with semantics in LDR completion, the end-to-end execution takes 5,855 ms. We observe that the main impact factor for the LDR completion process is the number of environment map outputs. Increasing generation output introduces a near-linear yet slow increase in the estimation latency. Although our default system configuration uses five outputs, the estimation latency is only about double that of using a single estimation output. Note that these long execution time could be masked by pipelining per-frame requests to a powerful edge server, if needed. However, with our on-device refinement, we have the feasibility to skip expensive generative inferences yet still meet the real-time and quality goals. Specifically, once the high-intensity map is generated, CleAR can generate visually coherent HDR environment maps in real-time by matching the LDR environment map to the current camera view’s color appearance. If the environmental lighting conditions do not change much, our design allows for the reuse of expensive generative results across frames without sacrificing the visual quality. To put CleAR’s latency in context, we note that a recent AR-specific lighting estimation system LitAR (Zhao et al., 2022a) needs 46.6/134.38ms to generate low/high quality near-field portion of the environment map. Another AR-specific system Xihe can achieve real-time lighting estimation (20.1ms) but can only produce low-fidelity spherical harmonics coefficients. In contrast, CleAR can deliver real-time HDR environment maps, which are much higher quality lighting information.
@@ 6.2. Data-Driven Evaluation
Figure 8. Visualizations of two experiment protocols for data-driven evaluation. For the three-sphere protocol (left), we use environment maps (middle) from the Laval dataset to render virtual spheres of different material properties. For the robustness testing protocol (right), we use the robustness testing dataset, consisting of environment maps from Laval with edited ambient light intensity and color temperature. As an example, we show an environment map (middle) with four variants with different light intensities and color temperatures. The results (right) of light intensity augmentation are marked as dark and bright, and the results of color temperature augmentation are marked as warm and cold.
@@
We quantitatively evaluate the lighting estimation quality of CleAR against three state-of-the-art lighting estimation models: DiffusionLight (Phongthawee et al., 2023), StyleLight (Wang et al., 2022), and OmniDreamer (Akimoto et al., 2022). Among these baselines, DiffusionLight is most similar to CleAR because it uses both diffusion models and a multi-output generation pipeline. We use the Laval test set and its augmented variants (§5.2) to evaluate the accuracy and robustness of lighting estimation methods using the following two protocols:
@@
  • The three-sphere evaluation protocol (Wang et al., 2022; Phongthawee et al., 2023) uses a proxy method to evaluate the HDR lighting estimation accuracy through pixel-wise error measurements on three rendered spheres with different material properties. Specifically, the used sphere materials are: matte, silver matte, and silver. Figure 8 shows a set of rendered virtual spheres used in this evaluation protocol. Each material represents a mixture of specific reflectance and roughness properties. Compared to direct environment map-wise comparison, this evaluation protocol helps us understand the impact of lighting estimation on object rendering quality.
@@
  • The robustness testing protocol is designed by us to test the generalizability and robustness of lighting estimation systems under different lighting conditions, specifically changes in light intensity and color temperature. In this protocol, we test lighting estimation systems on the Laval variants and compare estimated environment maps with the ground truth to calculate pixel-wise errors. The comparison is performed in the LDR image domain to capture lighting condition-related color appearance differences between environment map images.
@@
Evaluation Metrics. Following prior works (Wang et al., 2022; Akimoto et al., 2022; Phongthawee et al., 2023; Gardner et al., 2019), we use the FID score (Seitzer, 2020) to measure the diversity of generated environment maps. We use three image-based metrics—scale-invariant Root Mean Square Error (si-RMSE) (Eigen et al., 2014), Angular Error (LeGendre et al., 2019), and RMSE—to evaluate lighting estimation accuracy on rendered virtual spheres. We follow  (Zhao et al., 2022a) in using RMSE to evaluate the accuracy of environment maps. We follow  (Phongthawee et al., 2023) to map the 0.1st and 99.9th value percentiles to 0 and 1 when calculating RMSE.
6.2.2. End-to-End Visual Quality and Performance.
@@
In Figure 9, we show qualitative comparisons between CleAR and baseline methods in three scenes, representing the neutral, warm, and cool lighting conditions. Compared to other methods, CleAR allows more accurate overall color tones rendering and creates more diverse reflection details on the rendered virtual spheres. Figure 10 presents a quantitative comparison between CleAR and two SoTA methods (Phongthawee et al., 2023; Wang et al., 2022) that can output HDR environment maps in both rendering quality and estimation latency. Under the same environmental observations, CleAR achieves the lowest scale-invariant and normalized RMSE values across all three virtual sphere types. Noticeably, for the mirror sphere, the most challenging material of the three, CleAR achieves approximately $50\%$ reduction in the scale-invariant RMSE values. On the other hand, we observe that CleAR generates environment maps with slightly higher angular errors. Upon further inspection, we suspect that the diverse environment map image details potentially caused the higher angular error because the angular error metric is sensitive to pixel-wise differences. Furthermore, CleAR achieves the lowest lighting estimation latency by large margins. In particular, compared to DiffusionLight (Phongthawee et al., 2023), a diffusion model-based method, CleAR takes significantly less time (110X). This reduction in estimation latency is due to CleAR’s fewer model inferences and lower per-inference latency compared to DiffusionLight. This result suggests that CleAR can enable more responsive user experiences in mobile AR applications.
Figure 9. Qualitative comparisons. We show examples of lighting estimation results on three scenes and the corresponding rendered virtual sphere images. For each method, we show rendered virtual spheres (row top) and estimated environment maps (row bottom). Note, OmniDreamer can only output LDR environment maps.
a RMSE
b Si-RMSE
c Angular Error
d Latency
Figure 10. End-to-end qualitative comparison. We compare the end-to-end estimation accuracy and latency between CleAR and two state-of-the-art works (Wang et al., 2022; Phongthawee et al., 2023). We test estimation accuracy through the three-sphere evaluation protocol, which directly measures the visual quality impact of environment lighting on AR virtual object rendering. CleAR achieves the lowest si-RMSE and RMSE values, on all three testing object material types in the three-sphere evaluation protocol. However, CleAR has slightly higher angular errors on the mirror and matte spheres. Moreover, CleAR achieves this competitive accuracy with significantly lower end-to-end latency compared to the other generative model-based methods.
@@
Impact of Context Data Availability. We follow the setups in (Wang et al., 2022; Phongthawee et al., 2023) to use a single centered view of a 75 horizontal FoV camera as the default configuration for CleAR. This configuration represents a baseline performance of CleAR because our system can take in more AR context data during runtime. Table 2 shows CleAR’s performance under different configurations of camera FoVs, the number of input views, and the usage of semantic maps. Note that the ambient light data has been included in all tests in Table 2 because it is inseparable from our estimation pipeline design. The added environment observations through new views and increased FoVs can both improve the accuracy of lighting estimation. Additionally, full-scene semantic maps are important guiding information for high-accuracy lighting estimation. Full-scene semantic maps can often be obtained as floor maps or scene design layouts in real-world applications. This ablation study suggests that sharing scene semantics can greatly benefit lighting estimation accuracy with our system.
Table 2. Impact of context data. We evaluate CleAR’s estimation accuracy under different configurations based on the three-sphere protocol. The first row shows the default configuration of CleAR, marked in gray, which already outperforms DiffusionLight and StyleLight in both si-RMSE and RMSE. As more context data is used—larger FoV, environmental semantics, and more camera observations, CleAR can further improve the rendering quality.
Scale-invariant RMSE $\downarrow$ Angular Error $\downarrow$ RMSE $\downarrow$
Configuration Diffuse Matte Mirror Diffuse Matte Mirror Diffuse Matte Mirror
RGB 75 FoV 1 view 0.06 0.15 0.29 4.42 5.92 8.89 0.17 0.21 0.37
RGB+semantics 75 FoV 1 view 0.06 0.14 0.28 4.41 5.91 8.89 0.17 0.20 0.37
RGB+semantics 110 FoV 1 view 0.05 0.13 0.27 4.19 5.77 8.66 0.16 0.19 0.35
RGB+semantics 110 FoV 3 views 0.03 0.10 0.24 3.89 5.21 8.23 0.13 0.17 0.31
RGB+semantics 110 FoV 5 views 0.02 0.09 0.22 3.71 5.13 8.17 0.11 0.16 0.29
RGB+full semantics 110 FoV 3 views 0.02 0.07 0.20 3.59 4.96 8.09 0.10 0.15 0.27
@@
Estimation Robustness Evaluation. We test the lighting estimation robustness of CleAR under several environmental lighting conditions with the robustness testing protocol. We compare CleAR against two recent lighting estimation methods that are built on top of generative models: OmniDreamer (Akimoto et al., 2022) and StyleLight (Wang et al., 2022)  [^6] We were unable to complete the robustness testing for DiffusionLight (Phongthawee et al., 2023) within a reasonable time given its high inference cost.. Figure 11 shows the RMSE values calculated on the estimated LDR environment maps. We choose to compare the LDR environment maps, instead of the HDR environment maps, because lighting conditions mainly affect the accuracy of LDR environment map estimation. We observe that the accuracy of both OmniDreamer and StyleLight is affected by the changing lighting intensities and color temperatures. While CleAR, on the other hand, shows consistently lower estimation error rates under different lighting conditions. Particularly, these two models exhibit much higher lighting estimation errors on lower environmental lighting intensity and warmer color temperature. The observed error increases potentially are caused by the unbalanced training datasets used by these models.
6.2.3. System Ablation Study.
In this section, we evaluate the estimation accuracy of CleAR by comparing its performance across different system configurations and design choices.
@@
a LDR Completion Output Diversity
b Color Refinement
c Output Selection
@@
Figure 12. Evaluation of generative estimation refinement. (12a) Our method achieves the lowest FID scores, showing strong generation diversity. (12b) Our color appearance matching technique improves estimation accuracy. (12c) Our output selection method reduces RMSE and variance compared to baselines.
Analysis of LDR Environment Map Completion. For the completion of the LDR environment map, we specifically examine the pixel-wise errors and the overall diversity of the environment map content. The former assesses CleAR’s LDR lighting estimation accuracy, and the latter evaluates the generation richness of CleAR generative estimation pipeline. For CleAR, we use the LDR environment maps after applying the color appearance adaptation. Following (Akimoto et al., 2022; Wang et al., 2022), we adopt two methods for measuring the generation content diversity: (i) calculate the FID score on the full estimated LDR environment map, and (ii) converting the environment map into a cube map and calculate FID scores on each face without the top and bottom faces as these two faces containing little information. In Table 12a, we show that our LDR environment map completion model can output environment map images with greater diversity than state-of-the-art models. Our LDR completion model outperforms other models by 27% to 370% in the first calculation method and 4% to 344% in the second calculation method. This observation confirms the effectiveness of our large LDR environment map completion dataset. Next, we compare the accuracy of the LDR environment map completion. Under the same environment observations as (Wang et al., 2022; Phongthawee et al., 2023), our method achieves comparable results to DiffusionLight even though our LDR completion model is based on a smaller-sized pre-trained diffusion model.
@@
Analysis of Estimation Refinement. We evaluate the performance of our on-device estimation refinement components on the impacts of lighting estimation accuracy. In Figure 12b, we show that our color appearance matching technique improves the overall estimation accuracy by 31%. The estimation refinement technique allows our system to achieve high-quality estimation with limited estimation outputs. By default, our system only requires five generation outputs while DiffusionLight (Yang et al., 2023) requires more than 90 generation outputs. Reducing the required generation output will also reduce the estimation latency and computation resources requirements during deployment. Next, we evaluate the effectiveness of our generation output selection policy. In Figure 12c, we show that our color-matching technique can reduce the estimation error rate compared to other selection methods. We observe that, with our output selection policy and a total of five generation outputs, CleAR can reduce 15% of the estimation error compared to using a single generation output. Furthermore, when using five-generation outputs, our generation output selection component can reduce the average estimation error rate by 11% compared to random selection.
6.3. Perceptual User Study